Interpret the empirical formula: Na_2SO_4 Interpre
Answer #1

wavelengh of the radiation = 3.83 \AA = 3.83 \times 10^{-10} m

we know that energy of 1 photon

E = h\frac{c}{\lambda }

where h= Planck's constant = 6.626 x 10¯34 J s

c = speed of light =  3.00 x 108 m/s

\lambda = wavelength = 3.83 \times 10^{-10} m

therefore putting the values E =5.19 x 10¯16 J